# MMSI-Bench: A Benchmark for Multi-Image Spatial Intelligence

### v1

@@ 4.1 Evaluation Setup
Benchmark models. We conduct a comprehensive evaluation of a wide range of MLLMs, encompassing both proprietary and open-source families, diverse model scales, and recent architectural advances. For proprietary models, our benchmark includes o3 [46], GPT-4.5 [45], GPT-4.1 [44], GPT-4o[43], Gemini-2.5-Pro [19] (with and without “Thinking” mode), Claude-3.7-Sonnet [10] (with and without “Thinking” mode), Doubao-1.5-pro [51], and Seed1.5-VL [24]. On the open-source side, we systematically assess state-of-the-art models such as Qwen2.5-VL [11], LLaVA-OneVision [30], InternVL3 [67], InternVL2.5 [16], DeepSeek-VL2 [61], Llama-3.2-11B-Vision [41], and NVILA [37].
Baselines. For comparison, we include three baselines: Random Guessing, Human Performance, and Blind GPT-4o. Random Guessing reflects the expected accuracy under equal-probability selection from all options. Human performance is measured as the average accuracy of five independent evaluators not involved in data annotation, representing human-level spatial intelligence on this benchmark. Blind GPT-4o denotes the performance of GPT-4o without access to image inputs.
Implementation details. We report accuracy (%) using exact match between answers extracted from model outputs and ground-truth answers for our multiple-choice questions. If a model fails to generate an answer in the required format, we adopt the LLM-based fallback strategy from VLMEvalKit [21] to extract the intended response. For consistency, all models are evaluated with a temperature of zero and a maximum output length of 2,048 tokens.
@@ 4.2 Main Results
Table 2 presents the primary results of our evaluation, covering both proprietary and open-source models, as well as the baselines of Random Guessing, Human Performance, and Blind GPT-4o. We summarize our main findings as follows:
Current MLLMs struggle with multi-image spatial reasoning. Despite recent progress, even the most advanced MLLMs exhibit significant limitations in multi-image spatial reasoning. Most models achieve relatively low average scores, often only marginally outperforming random guessing, and are far from human-level performance. For example, the best-performing proprietary reasoning model, o3, achieves an average accuracy of only 41.0%, while human performance exceeds 97%, highlighting a substantial gap in spatial intelligence between current MLLMs and humans. Furthermore, we find that employing the "thinking" mode in models such as Gemini-2.5-Pro and Claude-3.7-Sonnet brings only limited improvements, suggesting that these models potentially lack strong spatial reasoning capabilities. Additionally, the accuracy of Blind GPT-4o is close to that of random guessing, indicating that our tasks indeed require genuine visual-spatial reasoning and cannot be solved by language priors or commonsense knowledge alone.
@@ 4.3 Limited Effect of Prompting Techniques on MMSI-Bench
Prompting techniques are widely adopted to improve the reasoning and problem-solving capabilities of large models. Building on these advances, we investigate whether such methods can similarly enhance the spatial reasoning abilities of MLLMs on MMSI-Bench. Specifically, we investigate both linguistic and visual prompting approaches (see supplementary material for further details):
  • Linguistic prompting: We adopt the widely used Zero-Shot Chain-of-Thought (CoT) method. Following prior work [59], we prepend “Let’s think step by step” to the prompts to encourage reasoning.
  • Visual prompting: For multi-image spatial understanding, a fundamental requirement is the ability to identify overlaps between images in order to establish relationships across them. However, MLLMs may struggle to identify overlaps between images. To explicitly provide this cue to the model, we employ the popular image matching method PATS [42] to establish sparse correspondences between pairs of images, which involves concatenating multiple images and drawing lines between corresponding points to explicitly indicate cross-image relationships. As a baseline, we use concatenated images without correspondence lines as visual inputs, and compare their results with those obtained using visual prompt.
@@
We conduct experiments using representative models, including GPT-4o, GPT-4.5, LLaVA-OneVision-72B, and Qwen2.5-VL-72B. As shown in Figure 5, the linguistic prompting approach (CoT) leads to a modest performance improvement only for GPT-4o, while resulting to performance degradation on the other models. This limited effectiveness of linguistic prompting may be attributed to the models’ inherent lack of spatial reasoning capabilities; simply encouraging "thinking" does not address their fundamental limitations in spatial understanding—a finding consistent with the observation that adopting a "thinking" mode in models like Gemini-2.5-Pro and Claude-3.7-Sonnet does not yield performance gains either. The particularly pronounced negative impact for LLaVA-OneVision may be due to its limited exposure to multimodal reasoning data during training, which makes it struggle with complex multimodal reasoning tasks.
@@ 5.1 Error Type Categorization
Grounding errors. The model fails to correctly identify or localize relevant objects or details within the image. For instance, it might misclassify an object or incorrectly determine the position of an object—errors that undermine its ability to anchor visual reasoning in the actual image content.
Overlap-matching and scene-reconstruction errors. The model fails to identify and match corresponding points that represent the same locations or objects in the real scene across different images, and struggles to implicitly reconstruct the underlying scene based on these cross-image relationships. For example, it may be unable to recognize that a tree appearing in both images is actually the same tree, leading to an incorrect reconstruction of the scene layout.
Situation-transformation reasoning errors. The model makes mistakes when reasoning about spatial directions relative to different reference objects, or when converting between relative directions (such as left and right) and absolute directions (such as north, south, east, and west), leading to incorrect situational inferences.
Spatial-logic errors. The model exhibits errors in spatial logical reasoning, such as (1) hallucinating spatial relationships that do not exist in the given context; (2) failing to correctly apply the transitivity of spatial relations—for example, given that A is east of B and B is east of C, the model incorrectly infers that A is west of C; and (3) making mistakes in identifying the correct reference object during motion reasoning, such as using another moving object as the reference point when determining an object’s movement status (whereas the correct approach would be to use a stationary object like the ground as the reference).
@@ 5.2 Automated Error Analysis
The manually annotated reasoning processes that we provide, in addition to facilitating quality assurance, have other potential uses, such as serving as training data to improve the spatial reasoning capabilities of MLLMs [15, 64]. However, as our work focuses on evaluation benchmarks, how to scalably collect such data for training remains an open question for future work. Here, we demonstrate another important application: enabling automated error analysis.
Specifically, we leverage the reasoning processes generated by GPT-4o on randomly sampled data, as described in Section 5.1. By providing GPT-4o with both the ground-truth answers and the human-annotated reasoning steps, we prompt it to simultaneously assess the correctness of the MLLMs’ reasoning process and, if incorrect, identify the key error type responsible. For comparison, we assess the MLLMs’ reasoning processes by identifying the most critical error type responsible for each incorrect case. Ultimately, this automated error analysis closely aligns with human evaluators’ judgments, with 71.8% of classifications matching when human-annotated reasoning steps are available. In contrast, when these annotated reasoning steps are not provided and only the ground-truth answers are given, only 53.6% of classifications match human evaluators’ judgments, indicating that correct answers alone are insufficient for effectively assisting the model in evaluating the reasoning process. This highlights the important role of high-quality human annotations in achieving reliable automated error analysis.
@@ 6 Conclusion
To summarize, we introduce MMSI-Bench, a challenging and comprehensive benchmark specifically designed to assess the multi-image spatial reasoning capability of MLLMs. Our evaluation across 34 state-of-the-art models reveals a substantial gap between current MLLMs and human-level spatial intelligence. Leveraging expert-annotated reasoning processes, we further propose an automated error analysis pipeline that enables scalable and systematic diagnosis of model failures, highlighting concrete directions for improving spatial reasoning in future models. We hope MMSI-Bench will serve as a valuable resource for the community and accelerate progress toward more spatially capable and robust multimodal AI systems. We discuss the limitations in the supplementary material.
@@ Appendix A Appendix Overview and Organization
    Limitations, Social Impact, and License & Access (Section B): This section outlines not only the limitations of our work, but also discusses its potential social implications, associated risks, and ethical considerations. In addition, it clarifies the licensing terms and access conditions for MMSI-Bench, ensuring responsible and transparent use of the dataset.
@@
    Comparison with Other Spatial Question-Answering Benchmarks (Section C): In this section, we compare MMSI-Bench with existing spatial question-answering benchmarks, highlighting the unique value of our benchmark.
@@
    Additional Implementation Details and Experimental Results (Section D): This section provides supplementary technical details regarding the implementation as well as more experimental results.
@@
    Details of the Data Curation Process (Section E): This section elaborates on the procedures and criteria used in curating the MMSI-Bench dataset, including data sourcing, annotation protocols, and quality control measures.
@@
    Representative MMSI-Bench Samples from Each Category (Full Version) (Section F): In this section, we present a full set of representative samples from each category in MMSI-Bench, without any truncation or simplification. This is intended to give the reader a complete view of the dataset’s diversity and coverage.
@@
    Additional Case Studies (Section G): We provide further case studies to illustrate the model’s reasoning process on MMSI-Bench, supplementing those already discussed in the main text.
@@
Appendix B Limitations, Social Impact, and License & Access
B.1 Limitation
In developing MMSI-Bench, we opted for manual annotation of question-answer pairs to ensure higher data quality and richer content. While this approach leads to more challenging and diverse benchmarks, it inevitably introduces scalability limitations, as expanding the dataset requires significant human effort.
Additionally, during the manual annotation process, we intentionally avoided rigid templates to encourage greater variety and naturalness in the dataset. Although this results in a richer and more realistic set of questions, it also makes it more challenging to precisely control the difficulty level across all items.
Despite these inherent trade-offs, we believe that our annotation strategy enables MMSI-Bench to better reflect the complexities of real-world spatial reasoning, and will serve as a valuable resource for advancing research in vision-language understanding.
B.2 Broader Impact
@@
In summary, while MMSI-Bench is intended to advance responsible and transparent research in spatial reasoning for vision-language models, we urge the community to remain vigilant regarding both its positive and negative societal impacts. We are committed to updating and improving MMSI-Bench in response to constructive feedback, and we welcome ongoing dialogue to help identify and mitigate emerging risks.
B.3 Ethics Statement
As this research exclusively employs publicly accessible pre-trained models, it does not present any ethical concerns. All procedures strictly conform to the guidelines established by the NeurIPS Code of Ethics.
B.4 Reproducibility Statement
Our experiments follow the baseline configurations established by prior evaluation benchmarks or the original testing protocols of each respective model. Detailed implementation information is provided in the Evaluation Setup and Appendix sections. We will release all data and code, along with thorough documentation, to facilitate faithful reproduction of our main experimental findings. All aspects of this research are fully compliant with the NeurIPS Reproducibility Requirements.
B.5 License and Access
The MMSI-Bench dataset is distributed under the Creative Commons Attribution 4.0 International License (CC BY 4.0). For all images included in the dataset, their original licensing terms [3, 7, 6, 5, 4, 1, 2, 8] remain fully in effect. These images, as well as their annotations, are available for direct download at https://huggingface.co/datasets/sihany/MMSI-Bench.

### v2

@@ 4.1 Evaluation Setup
We evaluate 37 open-source and proprietary MLLMs on MMSI-Bench. For consistency, all models are evaluated with a temperature of zero and a maximum output length of 2,048 tokens.
For comparison, we include three baselines: Random Guessing, Human Performance, and Blind GPT-4o. Random Guessing reflects the expected accuracy under equal-probability selection from all options. Human performance is measured as the average accuracy of five independent evaluators not involved in data annotation, representing human-level spatial intelligence on this benchmark. Blind GPT-4o denotes the performance of GPT-4o without access to image inputs.
We report accuracy (%) using exact match between answers extracted from model outputs and ground-truth answers for our multiple-choice questions. If a model fails to generate an answer in the required format, we adopt the LLM-based fallback strategy from VLMEvalKit (Duan et al., 2024) to extract the intended response. More experimental details are provided in Appendix E.
@@ 4.2 Main Results
Table 3 presents the primary results of our evaluation. We summarize our main findings as follows:
Current MLLMs struggle with multi-image spatial reasoning. Despite recent progress, even the most advanced MLLMs exhibit significant limitations in multi-image spatial reasoning. Most models achieve relatively low average scores, often only marginally outperforming random guessing, and are far from human-level performance. For example, the best-performing proprietary reasoning model, GPT-5, achieves an average accuracy of only 41.9%, while human performance exceeds 97%, highlighting a substantial gap in spatial intelligence between current MLLMs and humans. Additionally, the accuracy of Blind GPT-4o is close to that of random guessing, indicating that our tasks indeed require genuine visual-spatial reasoning and cannot be solved by language priors or commonsense knowledge alone.
@@
Table 3: Evaluation results for 37 MLLMs on MMSI-Bench. For each category, the best-performing proprietary model and the best-performing open-source model are both indicated in bold.
Models Positional Relationship Attribute Motion MSR Avg. Cam.–Cam. Obj.–Obj. Reg.–Reg. Cam.–Obj. Obj.–Reg. Cam.–Reg. Meas. Appr. Cam. Obj. Proprietary GPT-5 43.0 35.1 32.1 48.8 42.4 51.8 60.9 36.4 32.4 36.8 42.0 41.9 o3 45.2 39.4 37.0 44.2 47.1 62.6 54.7 28.8 31.1 32.9 34.9 41.0 GPT-4.5 34.4 29.8 39.5 51.2 47.1 55.4 39.1 33.3 41.9 40.8 36.4 40.3 GPT-4.1 36.6 26.6 27.2 29.1 36.5 27.7 37.5 24.2 36.5 32.9 28.8 30.9 GPT-4o 34.4 24.5 23.5 19.8 37.6 27.7 32.8 31.8 35.1 36.8 30.8 30.3 Gemini-2.5-Pro 39.7 31.9 39.5 45.3 35.2 43.3 51.5 21.2 36.4 30.2 34.3 36.9 Claude-3.7-Sonnet 35.5 29.8 29.6 31.4 32.9 34.9 37.5 25.8 23.0 31.6 25.8 30.2 dots.vlm 1 33.3 31.9 38.3 27.9 34.1 42.2 62.5 36.4 23.0 25.0 30.8 34.1 GLM-4.5V-thinking 38.7 35.1 29.6 30.2 40.0 39.8 45.3 28.8 25.7 32.9 30.3 33.8 Doubao-1.5-pro 25.8 33.0 33.3 33.7 49.4 39.8 37.5 31.8 27.0 26.3 29.8 33.0 Seed1.5-VL 32.2 30.8 25.9 23.2 38.8 32.5 39.0 21.2 36.4 25.0 26.2 29.7 Open-source InternVL3-78B 34.4 23.4 32.1 12.8 37.6 26.5 37.5 19.7 28.4 31.6 29.3 28.5 InternVL2.5-78B 23.7 22.3 39.5 29.1 31.8 42.2 35.9 19.7 17.6 26.3 27.3 28.5 Qwen2.5-VL-72B 25.8 34.0 34.6 23.3 34.1 36.1 45.3 27.3 27.0 30.3 27.3 30.7 LLaVA-OneVision-72B 43.0 31.9 33.3 30.2 37.6 38.6 28.1 19.7 13.5 32.9 15.7 28.4 InternVL3-38B 21.5 20.2 33.3 23.3 35.3 25.3 39.1 21.2 16.2 31.6 25.8 26.3 InternVL2.5-38B 18.3 22.3 35.8 22.1 38.8 34.9 37.5 25.8 14.9 38.2 25.3 27.9 Qwen2.5-VL-32B 24.7 26.6 29.6 22.1 32.9 31.3 31.2 24.2 18.9 35.5 27.8 27.7 InternVL2.5-26B 24.7 19.1 29.6 33.7 31.8 37.3 35.9 30.3 10.8 31.6 26.8 28.0 NVILA-15B 30.1 39.4 28.4 36.0 38.8 20.5 29.7 31.8 18.9 35.5 27.8 30.5 InternVL3-14B 19.4 24.5 24.7 23.3 37.6 24.1 31.2 22.7 24.3 31.6 29.3 26.8 Llama-3.2-11B-Vision 25.8 30.8 32.0 25.6 21.2 25.9 20.3 19.7 25.6 28.9 19.2 25.4 InternVL3-9B 18.3 25.5 32.1 29.1 31.8 22.9 29.7 24.2 16.2 38.2 26.8 26.7 InternVL3-8B 25.8 31.9 37.0 25.6 35.3 28.9 23.4 24.2 16.2 32.9 14.6 25.7 InternVL2.5-8B 32.3 27.7 29.6 32.6 24.7 32.5 26.6 27.3 16.2 31.6 30.3 28.7 NVILA-8B 17.2 29.8 24.7 30.2 22.4 34.9 34.4 25.8 25.7 34.2 29.8 28.1 Qwen2.5-VL-7B 24.7 24.5 24.7 25.6 29.4 26.5 25.0 18.2 20.3 39.5 25.8 25.9 LLaVA-OneVision-7B 20.4 33.0 29.6 29.1 25.9 30.1 29.7 25.8 18.9 34.2 11.6 24.5 InternVL2.5-4B 31.2 23.4 21.0 31.4 34.1 25.3 23.4 24.2 13.5 31.6 26.8 26.3 Qwen2.5-VL-3B 26.9 27.7 30.9 29.1 28.2 34.9 31.2 16.7 17.6 27.6 23.2 26.5 InternVL3-2B 26.9 25.5 29.6 31.4 28.2 27.7 26.6 22.7 12.2 23.7 23.7 25.3 InternVL2.5-2B 28.0 27.7 24.7 37.2 29.4 36.1 43.8 15.2 21.6 31.6 26.8 29.0 InternVL3-1B 24.7 35.1 22.2 30.2 29.4 30.1 32.8 28.8 17.6 19.7 26.3 27.0 InternVL2.5-1B 23.7 26.6 24.7 25.6 31.8 25.3 31.2 30.3 17.6 25.0 26.3 26.1 DeepSeek-VL2 23.7 31.9 22.2 36.0 30.6 22.9 28.1 15.2 28.4 26.3 28.3 27.1 DeepSeek-VL2-Small 24.7 28.7 18.5 33.7 38.8 27.7 28.1 33.3 24.3 25.0 29.8 28.6 DeepSeek-VL2-Tiny 29.0 27.7 21.0 23.3 17.6 31.3 14.1 24.2 14.9 25.0 27.3 24.0 Baseline Blind GPT-4o 20.2 17.0 29.6 13.9 29.4 19.2 21.8 12.1 20.2 29.0 20.2 22.7 Random Guessing 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 25.0 Human Level 95.7 98.9 97.5 94.2 98.8 96.4 95.3 98.5 98.6 98.7 97.0 97.2
Current spatial fine-tuning yields limited benefits on MMSI-Bench. We evaluated models fine-tuned on spatial-reasoning data on MMSI-Bench (detailed results are provided in Appendix E.4) and found only marginal gains over their base counterparts, underscoring the benchmark’s high difficulty and stringent demands for spatial reasoning generalization.
@@ 4.3 Limited Effect of Prompting Techniques on MMSI-Bench
Prompting techniques are widely adopted to improve the reasoning and problem-solving capabilities of large models. Building on these advances, we investigate whether such methods can similarly enhance the spatial reasoning abilities of MLLMs on MMSI-Bench. Specifically, we investigate both linguistic and visual prompting approaches (see Appendix E.2 for further details):
Linguistic prompting: We adopt the widely used Zero-Shot Chain-of-Thought (CoT) method. Following prior work (Wei et al., 2022), we prepend “Let’s think step by step” to the prompts to encourage reasoning.
Visual prompting: For multi-image spatial understanding, a fundamental requirement is the ability to identify overlaps between images in order to establish relationships across them. However, MLLMs may struggle to identify overlaps between images. To explicitly provide this cue to the model, we employ the popular image matching method PATS (Ni et al., 2023) to establish sparse correspondences between pairs of images, which involves concatenating multiple images and drawing lines between corresponding points to explicitly indicate cross-image relationships. As a baseline, we use concatenated images without correspondence lines as visual inputs, and compare their results with those obtained using visual prompt.
@@
We conduct experiments using 4 representative models. As shown in Figure 5, the linguistic prompting approach (CoT) leads to a modest performance improvement only for GPT-4o, while resulting in performance degradation on the other models. This limited effectiveness of linguistic prompting may be attributed to the models’ inherent lack of spatial reasoning capabilities; simply encouraging "thinking" does not address their fundamental limitations in spatial understanding.
@@ 5.1 Error Type Categorization
Grounding errors. The model fails to correctly identify or localize relevant objects or details within the image. For instance, it might misclassify an object or incorrectly determine the position of an object, errors that undermine its ability to anchor visual reasoning in the actual image content.
Overlap-matching and scene-reconstruction errors. The model fails to identify and match corresponding points that represent the same locations or objects in the real scene across different images, and struggles to implicitly reconstruct the underlying scene based on these cross-image relationships. For example, it cannot recognize that a tree appearing in both images is actually the same tree, leading to an incorrect reconstruction of the scene layout.
Situation-transformation reasoning errors. The model makes mistakes when reasoning about spatial directions relative to different reference objects, or when converting between relative directions (such as left and right) and absolute directions (such as north, south, east, and west).
Spatial-logic errors. The model exhibits errors in spatial logical reasoning, such as (1) hallucinating spatial relationships that do not exist in the given context; (2) failing to correctly apply the transitivity of spatial relations, for example, given that A is east of B and B is east of C, the model incorrectly infers that A is west of C; and (3) making mistakes in identifying the correct reference object during motion reasoning, such as using another moving object as the reference point when determining an object’s absolute motion status (whereas the correct approach would be to use a stationary object like the ground as the reference).
@@ 5.2 Automated Error Analysis
Error analysis of model outputs can provide valuable insights for improving MLLMs’ spatial capabilities. However, manually reviewing each output is time-consuming and impractical at scale.
We try to give the model the correct answer and ask it to reflect on its output and identify the primary error type following Section 5.1. In experiments with GPT-4o, it correctly labels the error type for only 53% of the selected samples, indicating that the correct answer alone is insufficient for evaluating the reasoning process. When we also supply the human-annotated reasoning process, the accuracy rises to 78%, highlighting the effectiveness and essential role of our reasoning annotation in achieving reliable automated error analysis.
@@ 6 Conclusion
To summarize, we introduce MMSI-Bench, a challenging and comprehensive benchmark specifically designed to assess the multi-image spatial reasoning capability of MLLMs. Our evaluation across 37 state-of-the-art models reveals a substantial gap between current MLLMs and human-level spatial intelligence. Leveraging expert-annotated reasoning processes, we further propose an automated error analysis pipeline that enables scalable and systematic diagnosis of model failures, highlighting concrete directions for improving spatial reasoning in future models. We hope MMSI-Bench will serve as a valuable resource for the community and accelerate progress toward more spatially capable and robust multimodal AI systems. We further discuss the LLM usage and limitations in Appendix B.
@@ Appendix A Appendix Overview and Organization
    LLM Usage, Limitations, Social Impact, and License & Access (Section B): This section outlines not only our LLM Usage, the limitations of our work, but also discusses its potential social implications, associated risks, and ethical considerations. In addition, it clarifies the licensing terms and access conditions for MMSI-Bench, ensuring responsible and transparent use of the dataset.
@@
    Motivation for Multi-Image Input (Section C): We discuss the unique advantages of multi-image input for evaluating spatial intelligence.
@@
    Comparison with Other Spatial Question-Answering Benchmarks (Section D): In this section, we compare MMSI-Bench with existing spatial question-answering benchmarks, highlighting the unique value of our benchmark.
@@
    Additional Implementation Details and Experimental Results (Section E): This section provides supplementary technical details regarding the implementation as well as more experimental results.
@@
    Details of the Data Curation Process (Section F): This section elaborates on the procedures and criteria used in curating the MMSI-Bench dataset, including data sourcing, annotation protocols, and quality control measures, as well as the methodology and results of the difficulty-level annotation.
@@
    Representative MMSI-Bench Samples from Each Category (Full Version) (Section G): In this section, we present a full set of representative samples from each category in MMSI-Bench, without any truncation or simplification. This is intended to give the reader a complete view of the dataset’s diversity and coverage.
  7. 7.
    Additional Case Studies (Section H): We provide further case studies to illustrate the model’s reasoning process on MMSI-Bench, supplementing those already discussed in the main text.
@@
Appendix B LLM Usage, Limitations, Social Impact, and License & Access
B.1 LLM Usage
In the preparation of this manuscript, Large Language Models (LLMs) were used as a writing-assistance tool. The primary purpose of their use was to aid in polishing the language, which included improving grammar, refining sentence structure, and enhancing the overall clarity and readability of the text. All core research concepts, methodologies, analyses, and conclusions were conceived and formulated exclusively by the authors. The authors have reviewed the final text and take full responsibility for the entire content of this paper, ensuring its scientific accuracy and originality.
B.2 Limitations
To ensure higher data quality and richer content, we chose to manually annotate question–answer pairs. This approach may limit scalability because it requires substantial human effort, which we acknowledge as a potential limitation of MMSI-Bench. Nevertheless, MMSI-Bench’s current scale (1,000 samples) is sufficient for meaningful assessment, given that today’s MLLMs still have substantial headroom for improvement on this benchmark. Its size is also comparable to widely used VQA benchmarks, such as MMStar (1,500 samples), MM-Vet (218 samples), and LLaVA-in-the-Wild (60 samples). When performance begins to saturate, a larger benchmark may indeed be necessary, but reaching that point will likely require new research breakthroughs or more human annotation effort.
Notably, MMSI-Bench’s scale is competitive among spatial benchmarks. For reference: ERQA has 400 samples, MuriBench 478, UniQA-3D 500, SpatialVLM 546, BLINK 572, LEGO-Puzzle 1,100, SpatialRGPT 1,406, and VSI-Bench 5,000. MMSI-Bench contains 1,000 samples, of which only the template-based VSI-Bench is substantially larger.
B.3 Broader Impact
@@
In summary, while MMSI-Bench is intended to advance responsible and transparent research in spatial reasoning for vision-language models, we urge the community to remain vigilant regarding both its positive and negative societal impacts.
B.4 Ethics Statement
As this research uses only publicly accessible pre-trained models, it does not present any ethical concerns. All procedures strictly conform to the guidelines established by the ICLR Code of Ethics.
B.5 Reproducibility Statement
Our experiments follow the baseline configurations established by prior evaluation benchmarks or the original testing protocols of each respective model. Detailed implementation information is provided in the Evaluation Setup and Appendix sections. We will release all data and code, along with thorough documentation, to facilitate faithful reproduction of our main experimental findings. All aspects of this research are fully compliant with the ICLR Reproducibility Requirements.
B.6 License and Access
The MMSI-Bench dataset is distributed under the Creative Commons Attribution 4.0 International License (CC BY 4.0). For all images included in the dataset, their original licensing terms (dtu, , sca, , nus, , hab, , ego, , agi, , dav, , way, ) remain fully in effect.